A sample of gas in a cylinder as in the example in Part A has an initial volume of 42.0 L , and you have determined that it contains 1.50 moles of gas. The next day you notice that some of the gas has leaked out. The pressure and temperature remain the same, but the volume has changed to 10.5 L . How many moles of gas (n2) remain in the cylinder?

Express your answer with the appropriate units.

View Available Hint(s)
Answer #1

According to Avogadro's law, at constant temperature and pressure volume of gas is directly proportional to the number of moles of gas present in it.

V​​​​​​1/V​​​​​​2 = n​​​1 /n​2  

42.0 L/10.5 L = 1.50 moles/n​​​​​​2

n​​​​​​2​​​ = 1.50×10.5/42.0 moles = 0.375 moles

Please give feedback!